# Normalize Filters! Classical Wisdom for Deep Vision

### v1

@@ 1 Introduction
$+$ extra blurring $=$unnormali zed filtering
@@
Filter normalization is critical because it prevents artifacts such as intensity shifts, halos, and ringing, which can distort the output and undermine the reliability of downstream tasks. For example, in Gaussian blurring, the filter kernel is normalized so that its weights sum to $1$, preserving the overall intensity of the image while smoothing out fine details. Similarly, Gaussian derivative filters are normalized so that its positive weights sum to $+1$ and negative weights sum to $-1$, enabling accurate edge detection and allowing edge strengths to be compared across different types of derivative filters. This careful normalization process has been a hallmark of classical filtering, ensuring consistency and robustness across a wide range of applications (See Fig. 1).
@@
In Fig. 1, for the checkerboard under uniform and uneven illumination, the normalized Difference of Gaussian (DoG) filter consistently detects checker edges. In contrast, the unnormalized DoG filter, whose positive weights sum to $+2$ and negative weights to $-1$, is equivalent to the sum of the normalized DoG and a normalized blur filter. This extra blurring causes the response to shift with illumination and overwhelm the edge response of the normalized DoG.
@@ 2 Related work
Deep networks’ performance can be severely impacted by atmospheric transformations. To mitigate this, various approaches have been developed such as architecture modifications to achieve equivariance to some transformations, normalization layers to enhance network robustness to instance-level perturbations, and data augmentations to improve robustness and generalization.
@@ 3 Filter normalization for deep learning
3.1 Any filter is both averaging and differencing
@@
$w^{-}$ $=(-w)\circ 1(w<0)\vspace{-1pt}$ (4)
[…]
Figure 4: Convolutional filters learned in CNNs are unnormalized, leaning more toward differencing than averaging. The distribution of the positive weight ratio of filters for each model optimized for ImageNet classification is skewed towards 00, indicating that most filters primarily perform differencing.
@@
    $r(w)\!=\!1$ ($\!-\!1$) when and only when all non-zero weights of $w$ are positive (negative).
@@
[…]
@@
We show that an unnormalized filter $w$ is a weighted sum of a differencing filter and an averaging filter (Fig. 1). Without loss of generality, we assume $\|w^{+}\|\geq\|w^{-}\|$; otherwise, we can study $-w$ instead. We have:
@@
Therefore, a general filter behaves between averaging and differencing, and the closer its positive weight ratio is to 00 ($1$), the more dominant differencing (averaging) it becomes.
@@ 3.2 Filter normalization
If $g\!>\!1$ ($g\!<\!1$), the image looks brighter (dimmer) with an expanded (reduced) range; if $g\!<\!1$ and $o\!>\!0$, the image looks brighter and hazy with reduced contrast (Fig. 2). For any unnormalized filter $w$, the response becomes
$f(g\cdot x+o)$ $=g\cdot f(x)+o\cdot\|w\|\cdot r(w)=\begin{cases}g\cdot f(x)+o,&\text{if }w\text{ is averaging}\\ g\cdot f(x),&\text{if }w\text{ is differencing}\end{cases}.$ (11)
That is, normalized filters produce consistent and interpretable responses with respect to intensity changes: Both averaging and differencing filters are equivariant to intensity gain modulation and shifts; in particular, differencing filters are invariant to intensity shifts.
If all filters in a layer are normalized, their responses are scaled uniformly by the same gain factor $g$. For unnormalized filters, however, their responses will be scaled differently, each by the filter’s total weight sum, $\|w\|r(w)$, thereby altering the feature representation at that layer.
@@
where $\varepsilon$ is a small constant (e.g., $10^{-6}$) to ensure numerical stability. This normalization step enforces the filter to become either averaging if $w$ is all positive, or differencing if $w$ has both positive and negative weights, ensuring atmosphere equivariance and enabling co-domain symmetry. It is then followed by learnable scaling $a$ and shifting $b$, which model response weighting $\|w\|$ and offset $b$.
@@ 4.1 Results on artificially corrupted benchmarks
We propose four additional evaluation sets by applying random gain and bias perturbations to $D$. 1) $D_{C}$ has constant random gain-bias corruptions to the entire image: $\tilde{x}=\mathbf{1}(\alpha x+\beta$), where $\tilde{x}$ is the corrupted image, and $\alpha\sim\text{Unif}(0.7,1.3)$ and $\beta\sim\text{Unif}(-0.3,0.3)$ are calculated to have a maximum variation of ±$\pm 30\%$ from the original image $x$ intensity. 2)$D_{L}$ has gain-bias corruption in a smooth linearly decreasing fashion: $\tilde{x}=x\cdot\textbf{L}_{\alpha}+\textbf{L}_{\beta}$, where $\textbf{L}_{\alpha}$ is a linearly varying field with values $[\alpha_{0},\alpha_{1}]$ where $\alpha_{0},\alpha_{1}\sim\text{Unif}(0.5,1.5)$ and $\textbf{L}_{\beta}$ a different linear field with values $[\beta_{0},\beta_{1}]$ where $\beta_{0},\beta_{1}\sim\text{Unif}(-0.5,0.5)$, both fields with the same random direction. 3) $D_{B}$ has the gain-bias perturbation in a blob with a fixed but decaying radius centered on a random pixel in the image: $\tilde{x}=x\cdot\textbf{B}_{\alpha}+\textbf{B}_{\beta}$, where $\textbf{B}_{\alpha}$ and $\textbf{B}_{\beta}$ are blob fields calculated with a cubic decay over a radius of 0.8 the size of the image. $D_{C},D_{L},D_{B}$ simulate global and local atmospheric effects such as variations in illumination caused by the environment or artificial light sources, calibration, or artifacts from sensor defects (See Fig. 2). 4) $D_{S}$ perturbation is also constant as in $D_{C}$, but with a fixed strong shifted bias to increase the data distribution shift: $\tilde{x}=\mathbf{1}(\alpha x+\beta+\gamma$), where $\alpha\sim\text{Unif}(0.7,1.3)$, $\beta\sim\text{Unif}(-0.3,0.3)$ and $\gamma=1$.
@@
We evaluate our proposed normalized convolutions on CIFAR with vanilla ResNets resnet . Additionally, we compare to norm-equivariant nets herbreteau2023 , which is the most recent baseline and closest related to our method. We replace the vanilla convolutions and activations in the ResNets with affine convolutions and sort pooling layers, respectively, from their official implementation.
We train all models using SGD with a learning rate of 0.1, a cosine annealing schedule for 200 epochs, and a batch size of 128. Fig. 5bottom shows our approach’s superior robustness to atmospheric effects, outperforming vanilla R20 and norm-equivariant R20E. See complete results with larger models in Table A1.
To further analyze the impact of corruptions on feature representations, we visualize the last layer features on CIFAR-10’s test set using t-SNE tsne . As shown in Fig. 6top, a vanilla R20 struggles to produce discriminative features in the presence of atmospheric variations, whereas our approach (Fig. 6bottom) preserves features more effectively even under severe corruptions. In Appendix A.1.1 we show the performance of R20 and R20N on $D_{C}$ varying the amount of corruption.
Figure 6: Feature quality from unnormalized convolutions drops under atmospheric perturbations. We show the t-SNE visualization of a R20 last layer’s features with moderate and strong levels of corruption on CIFAR-10. Top: A vanilla R20 fails under atmospheric perturbations, while Bottom: our approach maintains the feature quality even with strong perturbations.
@@
Figure 9: Our convolutions maintain their focus on salient regions despite the presence of atmospheric effects. Grad-CAM visualizations show that a vanilla R34 trained on ImageNet loses focus on relevant regions under corruptions (top), while our R34N produces more stable feature representations (bottom).
@@
Notably, our R34N with 22M parameters achieves better results than larger models like ResNet101 (45M), ViT-B (87M), and CLIP clip (151M) trained on much larger datasets like LAION laion .
@@ 4.3 Analysis and ablation studies
In Fig. 14right we show that for a R34, the images with a higher response from filters with higher $|r(w)|$ are more likely to be misclassified. For instance, filters with highest $|r(w)|$ misclassify $\sim$67of the images, while the least unnormalized misclassify $\sim$30%. Our R34N avoid this issue by design.
@@
To evaluate the effectiveness of data augmentation, we train a vanilla R20 and our proposed R20N using atmospheric augmentations on CIFAR-10. We use random gain-bias augmentations with a maximum variation of ±$\pm 10\%$ from the original image intensity to simulate the case where we want to approximate the testing data distribution but we do not know the exact distribution (e.g., ±$\pm 30\%$ maximum variation on $D_{C}$).
@@
While both IN and our approach mitigate instance-specific biases, ours offers two key distinctions: 1) independent normalization of weights, which reduces artifacts; and 2) normalization of weights rather than activations, providing a more targeted correction.
@@ Appendix A Supporting experiments
The supplementary material provides more comprehensive tables with results from our main experiments and ablations on CIFAR-10 in § 4. Specifically, tables providing numerical results from Figures 5, 8, 8, and 14 in § A.1A.2A.3, and A.4. We also provide an ablation with an increasing amount of corruption in § A.1.1, and analysis on enforcing normalization as a soft regularization in contrast to our method in § A.5. Additionally, we include the training curves of the vanilla R34 and our R34N trained on ImageNet-1k in § A.6 and learned filter visualizations in § A.7.
@@ A.5 Normalization as soft regularization
$\begin{split}R(W)=\sum_{w\in W}\Bigl{(}\bigl{|}1-\|w^{+}\|\bigl{|}+\bigl{|}1-\|w^{-}\|\bigl{|}\Bigl{)}.\end{split}$ (14)
@@
A.6 R34 training curves on ImageNet
@@ A.7 Feature visualizations
A.8 Results on ImageNet-C and ImageNet-P
We train a R34 and our R34N on the clean ImageNet-1k training set and evaluate them on the ImageNet-C and ImageNet-P datasets imagenetc . ImageNet-C tests model robustness by applying various corruptions to images, while ImageNet-P evaluates model stability by applying 30 levels of increasing perturbations and measuring the Flip Rate (FR), which is the frequency of changed predictions from the original image’s predicted class. For ImageNet-C, we assess performance across all corruption levels, and for ImageNet-P, we assess performance at the maximum perturbation level. We evaluate the weather and noise corruption groups for both datasets, as they are the most relevant for this work.
In Table A7, we observe comparable results, which we attribute to the clipping of ImageNet-C and ImageNet-P images are saved to disk. This clipping likely causes information loss, particularly under extreme corruptions like high gain or bias, resulting in overexposed regions with saturated pixel values (e.g., pure white). Unlike scientific imaging formats like FITS, which preserve dynamic range, the clipped images may limit the effectiveness of our approach.
Table A7: Results on ImageNet-C and ImageNet-P. We show the Top-1 accuracy (Acc. $\uparrow$) for ImageNet-C and the Flip Rate (FR $\downarrow$) for ImageNet-P, focusing on weather and noise corruption groups. We observe comparable results, likely due to clipping that occurs when ImageNet-C and ImageNet-P images are saved to disk.
ImageNet-C (Acc. $\uparrow$) ImageNet-P (FR $\downarrow$) Model Weather Noise Weather Noise R34 41.6 27.4 41.3 23.5 R34N 41.0 28.1 40.0 24.9

### v2

@@ 1 Introduction
++ extra blurring ==unnormali zed filtering
@@
Filter normalization is critical because it prevents artifacts such as intensity shifts, halos, and ringing, which can distort the output and undermine the reliability of downstream tasks. For example, in Gaussian blurring, the filter kernel is normalized so that its weights sum to $1$, preserving the overall intensity of the image while smoothing out fine details. Similarly, Gaussian derivative filters are normalized so that its positive weights sum to +1+1 and negative weights sum to 1-1, enabling accurate edge detection and allowing edge strengths to be compared across different types of derivative filters. This careful normalization process has been a hallmark of classical filtering, ensuring consistency and robustness across a wide range of applications (See Fig. 1).
@@
In Fig. 1, for the checkerboard under uniform and uneven illumination, the normalized Difference of Gaussian (DoG) filter consistently detects checker edges. In contrast, the unnormalized DoG filter, whose positive weights sum to +2+2 and negative weights to 1-1, is equivalent to the sum of the normalized DoG and a normalized blur filter. This extra blurring causes the response to shift with illumination and overwhelm the edge response of the normalized DoG.
@@ 2 Related work
Deep networks’ performance can be severely impacted by atmospheric transformations (we use“atmospheric” in the context of “The net effect of the viewing conditions, including additive and multiplicative effects” Adelson1999 ). To mitigate this, various approaches have been developed such as architecture modifications to achieve equivariance to some transformations, normalization layers to enhance network robustness to instance-level perturbations, and data augmentations to improve robustness and generalization.
@@ 3 Filter normalization for deep learning
3.1 Any linear filter is both averaging and differencing
@@
$w^{-}$ $=(-w)\circ 1(w<0)\vskip-1.0pt$ (4)
where both $w^{+}$ and $w^{-}$ are nonnegative $k\!\times\!1$ vectors. The $L_{1}$ norm $\|\cdot\|_{1}$ of these vectors is simply the total sum of their respective weights. We will use $L_{1}$ norm exclusively throughout this paper, with $\|\cdot\|_{1}$ as $\|\cdot\|$ for short.
Figure 4: Convolutional filters learned in CNNs are unnormalized, leaning more toward differencing than averaging. The distribution of the positive weight ratio of filters for each model optimized for ImageNet classification is skewed towards 0, indicating that most filters primarily perform differencing.
@@
    r(w)=1r(w)\!=\!1 (1\!-\!1) when and only when all non-zero weights of $w$ are positive (negative).
@@
    $r(w)\,\begin{subarray}{c}>\\ =\\ <\end{subarray}\,0$ when and only when $\|w^{+}\|\,\begin{subarray}{c}>\\ =\\ <\end{subarray}\,\|w^{-}\|$.
@@
We show that an unnormalized filter $w$ is a weighted sum of a differencing filter and an averaging filter (Fig. 1). Without loss of generality, we assume $\|w^{+}\|\geq\|w^{-}\|$; otherwise, we can study w-w instead. We have:
@@
Therefore, a general filter behaves between averaging and differencing, and the closer its positive weight ratio is to 0 (11), the more dominant differencing (averaging) it becomes.
@@ 3.2 Filter normalization
If g>1g\!>\!1 (g<1g\!<\!1), the image looks brighter (dimmer) with an expanded (reduced) range; if g<1g\!<\!1 and o>0o\!>\!0, the image looks brighter and hazy with reduced contrast (Fig. 2). For any unnormalized filter $w$, the response becomes
$\hskip-7.0ptf(g\cdot x+o)$ $=g\cdot f(x)+o\cdot\|w\|\cdot r(w).$ (11)
If all filters in a layer are unnormalized, each response is scaled by the filter’s total weight sum, $\|w\|r(w)$, thereby altering the feature representation at that layer.
@@
where $\varepsilon$ is a small constant (e.g., $10^{-6}$) to ensure numerical stability, and it is then followed by learnable scaling $a$ and shifting $b$, which model response weighting $\|w\|$ and offset $b$, respectively.
This normalization step enforces the filter to become either averaging if $w$ is all positive, or differencing if $w$ has both positive and negative weights, with predictable responses to intensity changes:
  1. 1.
    Averaging filters respond linearly to both gain and offset: $f(g\cdot x+o)=g\cdot f(x)+o$, where $\|w\|=1$ and $r(w)=1$.
  2. 2.
    Differencing filters are invariant to offset and equivariant to gain: $f(g\cdot x+o)=g\cdot f(x)$, where $\|w\|=1$ and $r(w)=0$.
This decomposition ensures atmosphere-equivariance and enable co-domain symmetry, meaning the filter response transforms in a predictable, structured way under affine transformations of the input intensity $x$, via gain $g$ and offset $o$, (i.e., global illumination changes.) Averaging filters respond to both contrast and brightness, while differencing filters respond only to contrast.
@@ 4.1 Results on artificially corrupted benchmarks
[…]
@@
We evaluate our normalized convolutions (R20N) on CIFAR with vanilla ResNets resnet (R20). Additionally, we compare to norm-equivariant nets herbreteau2023 (R20E), which is the most recent baseline and closest related to our method. Norm-equivariant networks replace ordinary convolutional layers with affine-constrained convolutions and traditional activation functions like ReLU with channel-wise sort pooling layers to preserve normalization-equivariance by design, allowing the network to handle changes in input scale and shift. Following their official implementation herbreteau2023 , we replace the vanilla convolutions and activations in the ResNets with affine convolutions and sort pooling layers, respectively.
Figure 6: Feature quality from unnormalized convolutions drops under atmospheric perturbations. We show the t-SNE visualization of a R20 last layer’s features with moderate and strong levels of corruption on CIFAR-10. Top: A vanilla R20 fails under atmospheric perturbations ($D_{c}$), while Bottom: our approach maintains the feature quality even with strong perturbations ($D_{C}$ with maximum variation of $\pm 100\%$).
We train the three models (R20, R20E, R20N) using SGD with a learning rate of 0.1, a cosine annealing schedule for 200 epochs, and a batch size of 128. Fig. 5bottom shows our approach’s superior robustness to atmospheric effects, outperforming vanilla R20 and norm-equivariant R20E. See complete results with larger models in Table A1.
To further analyze the impact of corruptions on feature representations, we visualize the last layer features on CIFAR-10’s test set using t-SNE tsne . As shown in Fig. 6top, a vanilla R20 struggles to produce discriminative features in the presence of atmospheric variations ($D_{C}$ ), whereas our approach (Fig. 6bottom) preserves features more effectively even under severe corruptions ($D_{C}$ with max. variation of $\pm 100\%$). In Appendix A.1.1 we show the performance of R20 and R20N on $D_{C}$ varying the amount of corruption.
@@
Figure 9: Our convolutions maintain their focus on salient regions despite the presence of atmospheric effects. Grad-CAM visualizations show that a vanilla R34 trained on ImageNet loses focus on relevant regions under strong corruptions (top), while our R34N produces more stable feature representations (bottom).
@@
Notably, our R34N with 22M parameters achieves better results than larger models like ResNet101 (45M), ViT-B (87M), and CLIP clip (151M) trained on much larger datasets like LAION laion . Although CLIP is used here as a zero-shot classifier and our R34N is trained specifically for ImageNet classification, the comparison is fair and meaningful because both models face the same unseen corruptions at test time. In fact, the comparison disadvantages our method, as CLIP was trained on 400M diverse images and likely encountered many degradations that our model has not seen. Despite this, R34N outperforms CLIP on intensity corruptions, highlighting the effectiveness of our filter normalization approach. On the original test set $D$, specialist models like R34N and R34 outperform generalist models like CLIP-B ($\sim$73% vs. 60.2%), reflecting their domain-specific focus. However, on corrupted sets, specialist baselines collapse (e.g., R34 drops from 73.3% to 2.1%), while CLIP-B drops but to a higher accuracy (from 60% to 14.3%). In contrast, R34N achieves 67.0% on corrupted sets, far surpassing both R34 and CLIP-B. This demonstrates that the robustness gain comes from the architecture itself, not from exposure to corruptions or tradeoffs on original clean performance.
@@ 4.3 Analysis and ablation studies
In Fig. 14right we show that for a R34, the images with a higher response from filters with higher $|r(w)|$ are more likely to be misclassified. For instance, filters with highest $|r(w)|$ misclassify $\sim$67% of the images, while the least unnormalized misclassify $\sim$30%. Our R34N avoid this issue by design.
@@
To evaluate the effectiveness of data augmentation, we train a vanilla R20 and our proposed R20N using atmospheric augmentations on CIFAR-10. We use random gain-bias augmentations with a maximum variation of $\pm 10\%$ from the original image intensity to simulate the case where we want to approximate the testing data distribution but we do not know the exact distribution (e.g., $\pm 30\%$ maximum variation on $D_{C}$).
@@
While both IN and our approach mitigate instance-specific biases, ours offers two key distinctions: 1) independent normalization of weights, which reduces artifacts that can occur when instance statistics are not representative of atmospheric corruptions; and 2) normalization of weights rather than activations, providing a more targeted correction by directly correcting the filter’s response to atmospheric corruptions, rather than indirectly adjusting the activations.
@@
Acknowledgments
This project was supported, in part, by NSF 2215542, NSF 2313151, and Bosch gift funds to S. Yu at UC Berkeley and the University of Michigan.
@@ A.5 Normalization as soft regularization
[…]
@@
A.6 Non-linearity impacts
Our proposed method ensures atmosphere-equivariance at the convolutional layer level. That is, prior to any non-linear activation, filter responses transform in a structured and predictable way under global intensity shifts (gain and bias). This forms the foundation for building robustness to such transformations throughout the network.
It is important to clarify that, while individual layers are equivariant, the goal of the full network is to achieve invariance to atmospheric transformations—just as classical CNNs preserve spatial equivariance locally but aim for global translational invariance in their final semantic outputs. Non-linearities such as ReLU serve this purpose: they progressively reduce sensitivity to irrelevant variations (e.g., illumination), transforming equivariant features into invariant representations through depth and learned hierarchy.
What makes our approach effective is that starting from an atmosphere-equivariant basis makes the job of learning invariance easier and more structured. This is evident in our experiments (Table 2), where models using filter normalization without any data augmentation outperform baselines with augmentation, by a large margin (e.g., +14% accuracy on $D_{S}$). This suggests that the network is better able to learn invariances downstream when the early layers encode meaningful, interpretable transformations.
In short, while non-linearities do break strict equivariance, they play a necessary and constructive role in turning structured responses into invariant representations. Our method complements this process by providing a more robust and interpretable foundation at the convolutional level.
We did further experiments on different types of activation functions, specifically Tanh and Sigmoid, in place of ReLU. Our results on CIFAR-10 $D_{C}$ with R20 and our R20N model show that our method consistently outperforms the vanilla R20 model, regardless of the activation function used (see Table A7).
Table A7: Filter normalization provides robustness benefits that are complementary to the choice of activation function, as demonstrated by experiments.
Model Tanh Sigmoid R34 90.5% 88.1% R34N 91.5% 91.5%
This suggests that our filter normalization approach provides robustness benefits that are complementary to the choice of activation function.
A.7 Further experimental validation
Our method is specifically designed to address atmospheric transformations—i.e., global additive and multiplicative corruptions such as changes in brightness, contrast, and haze (see Eq. 10 and Fig. 2). We already evaluate our method across a range of corruption types and intensities (see also Fig. A.1.1).
ImageNet-C imagenetc includes severe and often localized distortions (e.g., noise, blur, digital artifacts) that fall outside the atmospheric transformation family we explicitly target. Our method outperforms on some categories such as brightness and underperforms on others such as snow, all with a very small margin. While we acknowledge these results, we emphasize that ImageNet-P, proposed in the same paper, offers a more direct test of our model’s design goals: ImageNet-P applies smooth, progressive perturbations, especially in the "weather" category (e.g., brightness and snow), aligning well with our focus on global intensity transformations.
Robustness is measured using the Flip Rate (FR)—the fraction of samples whose predicted class changes at the highest perturbation level compared to the clean version: $FR=\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}(f(x_{i})\neq f(x^{30}_{i}))$, where $x^{30}_{i}$ is the corrupted image at the maximum perturbation level (30), and $n$ are all images included in the ”weather” category, which includes brightness and snow modifications (See Table A8).
The flip rate of a vanilla R34 is 41.3%, while our R34N with normalized convolutions achieves 40.0% flip rate (lower is better).
Table A8: Results on ImageNet-P. We show the Top-1 accuracy ($\uparrow$) and the Flip Rate (FR $\downarrow$) for ImageNet-P, focusing on weather corruptions (i.e., brightness and snow). The flip rate of a vanilla R34 is 41.3%, while our R34N with normalized convolutions achieves 40.0% flip rate (lower is better). While both metrics show consistent trends, the FR gap is more pronounced than the accuracy gap, making it a more sensitive indicator of perturbation robustness.
Model Accuracy ($\uparrow$) Flip Rate (FR $\downarrow$) R34 44.7 41.3 R34N 44.9 40.0
Lastly, we further assess generalization to natural variation by binning the ImageNet-1k test set into 9 contrast levels (based on pixel standard deviation). As shown in Table A9, R34N maintains more stable accuracy across the contrast spectrum compared to the baseline, which drops sharply at extreme values. This confirms the model’s improved robustness to natural realistic atmosphere conditions.
Table A9: Accuracy per contrast level. We bin the ImageNet-1k test set into 9 contrast levels. R34N maintains more stable accuracy across the contrast spectrum compared to the baseline, which drops sharply at extreme values.
Model Lowest Contrast 2 Contrast 3 Contrast 4 Contrast 5 Contrast 6 Contrast 7 Contrast 8 Highest R34 65.4% 74.0% 74.8% 75.2% 73.2% 70.8% 70.7% 67.9% 63.3% R34N 69.2% 74.1% 74.6% 74.5% 73.3% 71.1% 71.1% 67.3% 71.4%
Appendix B R34 training curves on ImageNet